OpenStudy (anonymous):

tell wnether the sequence is aritmetic. If the sequence is aritmetic, write a funcion rule to represent it 12, -9, -3...

OpenStudy (anonymous):

tell whether the sequence is aritmetic. If the sequence is aritmetic, write a funcion rule to represent it 12, -9, 6, -3... A no B A(n)= -3+(n-1)(-2) C A(n)= -3 + (n-1)(2) D A(n= 3-(n-1)(2)